# Influence of initial states on memory effects: A study of early-time superradiance

### v1

@@ IV.2 Factorized identical states
In this subsection, we investigate $N$-atom initial states in a factorized from
$\rho_{S}^{fact}(0)$ $=$ $\rho_{S_{1}}\otimes\rho_{S_{2}}\otimes...\otimes\rho_{S_{N}}$ (43)
@@
$\rho_{S_{1}}=\rho_{S_{2}}=...=\rho_{S_{N}}=\left(\begin{array}[]{cc}\rho_{ee}&\rho_{eg}\\ \rho_{eg}^{*}&\rho_{gg}\\ \end{array}\right).$ (46)
@@
Such states are fully characterized by $\rho_{ee}$ and $\rho_{eg}$ and may also be superradiant. For example, states of this form are experimentally studied in [13] to realize the single-atom superradiance where the a serious of atoms enter a cavity one by one. In [21], we show that such a model can be used to achieve the Heisenberg Limit in a quantum metrology scheme for measuring the atom-cavity coupling. The superradiance of two artificial atoms with initial state $\frac{1}{2}(|e\rangle+|g\rangle)\otimes(|e\rangle+|g\rangle)$ are experimentally observed in [15]. In the above works, the single-atom coherence ($|\rho_{eg}|$) are crucial for creating superradiance. Thus it is desirable to know the influence of such states on the memory effects in superradiance, especially the role of the single-atom coherence.
@@
In this subsection, the normalized value of memory effects and the normalized cavity photon number for states described by Eq.(43) are calculated using Eq.(27) and Eq.(30). When Eq.(46) is a pure state, i.e., $\rho_{Sn}=|\phi\rangle\langle\phi|$, $\rho_{S}^{fact}(0)$ can be decomposed into the superpositions of different Dicke states [13]. Using the analytical result of $\langle\phi|\sigma^{+}\sigma^{-}|\phi\rangle$ in [13], we obtain the normalized cavity photon number for $\rho_{S}^{fact}(0)$:
$\frac{N_{P}[\rho_{S}^{fact}(0)]}{(gt)^{2}}=N(N-1)|\rho_{eg}|^{2}+N\rho_{ee}.$ (47)
According to Eq.(32), the degree of superradiance is given by
$S[\rho_{S}^{fact}(0)]$ $=$ $\frac{N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}}{\sum_{n=1}^{N}N_{P}^{(n)}(\rho_{S_{n}})/(gt)^{2}}$
@@
$=$ $1+(N-1)|\rho_{eg}|^{2}/\rho_{ee}$
for $\rho_{ee}>0$. The result demonstrates that the state $\rho_{S}^{fact}(0)$ is superradiant when $|\rho_{eg}|$ is nonzero and $N>1$. Besides, the normalized cavity photon number as well as the degree of superradiance grow quadratically with the single-atom coherence $|\rho_{eg}|$ for $N>1$.
Figure 4: Normalized memory effects $N_{M}$ (left column) and normalized cavity photon number $N_{p}$ (right column) for $\rho_{S}^{fact}(0)$ as a function of $\rho_{ee}$ and $|\rho_{eg}|$ in the early-time regime. The atom numbers are $N=1$, $2$ and $11$ from top to bottom. For any $\rho_{ee}$ and $N$ ($N>1$), $N_{M}$ ($N_{p}$) can always be enhanced nonlinearly by the single-atom coherence $|\rho_{eg}|$.
In contrast to Eq.(40), the analytical solution of Eq.(27) for the factorized initial states is complicated in general. Therefore, Eq.(27) is solved numerically in this section. The results of the normalized $N_{M}[\rho_{S}^{fact}(0)]$ and $N_{P}[\rho_{S}^{fact}(0)]$ are depicted in Fig.4 as a function of $\rho_{ee}$ and $|\rho_{eg}|$ with $N=1$, 2 and $11$. The condition $|\rho_{eg}|^{2}\leqslant(1-\rho_{ee})\rho_{ee}$ are satisfied in Fig.4 to guarantee the positivity of $\rho_{S_{n}}$. The single-atom coherence $\rho_{eg}$ are chosen to be real in the simulations for simplicity. Numerical results shows that the phase of $\rho_{eg}$ does not affect the value of memory effects. Unlike the case for the Dicke states, $N_{M}[\rho_{S}^{fact}(0)]/(gt)^{2}$ is not proportional to $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ in general. However, our simulation results show that the normalized value of memory effects can always be enhanced by $|\rho_{eg}|$ for any $N$. Similar situations happen for $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ and $S[\rho_{S}^{fact}(0)]$ when $N>1$ as seen in Eq.(47) and (IV.2). Besides, when the atom number $N$ increases, $N_{M}[\rho_{S}^{fact}(0)]/(gt)^{2}$ grows more and more fast with $|\rho_{eg}|$, and so do $N_{P}[\rho_{S}^{fact}(0)]/(gt)^{2}$ and $S[\rho_{S}^{fact}(0)]$ as seen in Eq.(47) and (IV.2). Thus we conclude that the creation and enhancement of superradiance by $|\rho_{eg}|$ is accompanied by the increase of memory effects for the factorized identical initial states. In Fig.4(c) and (e), it is found that the states around $\rho_{eg}=0$ and $\rho_{ee}=0.5$ tend to take lower values of memory effects compared with their neighboring states, especially when $N$ is large. This might be intuitively interpreted by the fact that $\rho_{eg}=0$ and $\rho_{ee}=0.5$ leads to a maximally mixed state of the $N$-atom system. Besides, Fig.4 shows that the (change of) the cavity photon number is not the only factor to determine the value of the memory effects. Other factors might include the (change of) coherence of the cavity state between the basis $|n\rangle$.
@@ IV.3 Other initial states
Entanglement and coherence of the initial states may play important roles in superradiance [42, 22, 23, 24, 25], especially for the Dicke states. Next, we examine the influence of the coherence (or the entanglement among the atoms) in the dephased Dicke states $\rho^{deph}_{S}(0)$ on the memory effects as well as the superradiance. The initial state we consider is given by
$\rho^{deph}_{S}(0)=\lambda|JM\rangle\langle JM|+(1-\lambda)\mathcal{D}(|JM\rangle\langle JM|)$ (48)
where $0\leqslant\lambda\leqslant 1$. Here $\mathcal{D}$ turns $|JM\rangle\langle JM|$ into a fully dephased state with vanishing off-diagonal elements in the basis $|e(g),e(g),...,e(g)\rangle$. Meanwhile, $\mathcal{D}(|JM\rangle\langle JM|)$ is a mixed state of separable states whose entanglement is zero. Thus $\lambda$ reflects the strength of entanglement or the value of coherence of $\rho^{deph}_{S}(0)$. With the help of numerical solutions, we find by mathematical induction that $\frac{N_{M}[\mathcal{D}(|JM\rangle\langle JM|)]}{(gt)^{2}}=\frac{1}{2}(J+M)$ and
$\frac{N_{M}[\rho^{deph}_{S}(0)]}{(gt)^{2}}$ $=$ $\lambda\frac{N_{M}(|JM\rangle\langle JM|)}{(gt)^{2}}$ (49)
@@
$\frac{N_{P}[\rho^{deph}_{S}(0)]}{(gt)^{2}}=$ $=$ $\lambda\frac{N_{P}(|JM\rangle)}{(gt)^{2}}$ (50)
@@
Similarly, $N_{M}[\rho^{deph}_{S}(0)]=\frac{1}{2}N_{P}[\rho^{deph}_{S}(0)]$ is satisfied for the dephased Dicke states in the early-time regime. The degree of superradiance for the dephased Dicke states is given by
$S[\rho^{deph}_{S}(0)]$ $=$ $\frac{N_{P}[\rho^{deph}_{S}(0)]/(gt)^{2}}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)/(gt)^{2}}$ (51)
@@
for $J+M>0$. It is observed that the entanglement or coherence (represented by $\lambda$) in the depahsed Dicke state is necessary for superradiance. Besides, the degree of superradiance increases linearly with $\lambda$ for superradiant Dicke states. Similarly, the creation and increase of superradiance induced by $\lambda$ is accompanied by the increase of memory effects.
At the end of this subsection, we calculate the value of the normalized value of memory effects of a probabilistic mixture of Dicke states $\rho_{S}^{mix}(0)=\sum_{M}p_{M}|JM\rangle\langle JM|$ where $J=N/2$, $M=-J,-J+1,...,J$ and $\sum_{M}p_{M}=1$. Using similar treatment as in Eq.(39), we find that
$\frac{N_{M}[\rho_{S}^{mix}(0)]}{(gt)^{2}}\!$ $=$ $\!\frac{1}{4}[\!\sum_{M=-J}^{J-1}\!|p_{M+1}f(M+1)-p_{M}f(M)|$ (52)

### v2

@@ IV.2 Factorized identical states
In this section, we investigate $N$-atom initial states in a factorized from
$\rho_{S}^{\textrm{fact}}(0)$ $=$ $\rho_{S_{1}}\otimes\rho_{S_{2}}\otimes\cdots\otimes\rho_{S_{N}}$ (50)
@@
$\rho_{S_{1}}=\rho_{S_{2}}=\cdots=\rho_{S_{N}}=\left(\begin{array}[]{cc}\rho_{ee}&\rho_{eg}\\ \rho_{eg}^{*}&\rho_{gg}\\ \end{array}\right).$ (53)
@@
Such states are fully characterized by $\rho_{ee}$ and $\rho_{eg}$ and may also be superradiant. For example, states of this form are experimentally studied in Ref. [15] to realize the single-atom superradiance where a serious of atoms enter a cavity one by one. In Ref. [23], we show that such a model can be used to achieve the Heisenberg Limit in a quantum metrology scheme for measuring the atom-cavity coupling. The superradiance of two artificial atoms with initial state $\frac{1}{2}(|e\rangle+|g\rangle)\otimes(|e\rangle+|g\rangle)$ are experimentally observed in Ref. [17]. In the above works, the single-atom coherence ($|\rho_{eg}|$) are crucial for creating superradiance. Thus it is desirable to know the influence of such states on the memory effects in superradiance, especially the role of the single-atom coherence.
Figure 6: Normalized memory effects $N_{M}$ (left column) and normalized cavity photon number $N_{P}$ (right column) for $\rho_{S}^{\textrm{fact}}(0)$ as functions of $\rho_{ee}$ and $|\rho_{eg}|$ in the early-time regime. The atom numbers are $N=1$, $2$, and $11$ from top to bottom. For any $\rho_{ee}$ and $N$ ($N>1$), $N_{M}$ ($N_{P}$) can always be enhanced nonlinearly by the single-atom coherence $|\rho_{eg}|$.
@@
In the following, the normalized value of memory effects and the normalized cavity photon number for states described by Eq. (50) are calculated using Eqs. (29) and (33). When Eq. (53) is a pure state, i.e., $\rho_{Sn}=|\phi\rangle\langle\phi|$, $\rho_{S}^{\textrm{fact}}(0)$ can be decomposed into the superpositions of different Dicke states [15]. Using the analytical result of $\langle\phi|\sigma^{+}\sigma^{-}|\phi\rangle$ in Ref. [15], we obtain the normalized cavity photon number for $\rho_{S}^{\textrm{fact}}(0)$:
$\frac{N_{P}[\rho_{S}^{\textrm{fact}}(0)]}{(gt)^{2}}=N(N-1)|\rho_{eg}|^{2}+N\rho_{ee}.$ (54)
In contrast with Eq. (46), the analytical solution of Eq. (29) for the factorized initial states is more complicated in general. Therefore, Eq. (29) is solved numerically here. The results of the normalized $N_{M}[\rho_{S}^{\textrm{fact}}(0)]$ and $N_{P}[\rho_{S}^{\textrm{fact}}(0)]$ are depicted in Fig. 6 as functions of $\rho_{ee}$ and $|\rho_{eg}|$ with $N=1$, 2 and $11$. The condition $|\rho_{eg}|^{2}\leqslant(1-\rho_{ee})\rho_{ee}$ are satisfied in Fig. 6 to guarantee the semipositivity of $\rho_{S_{n}}$. The single-atom coherence $\rho_{eg}$ are chosen to be real in the simulations for simplicity. Numerical results shows that the phase of $\rho_{eg}$ does not affect the value of memory effects. Unlike the case for the Dicke states, $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/(gt)^{2}$ is not proportional to $N_{P}[\rho_{S}^{\textrm{fact}}(0)]/(gt)^{2}$ in general. It is observed that the normalized value of memory effects can always be enhanced by $|\rho_{eg}|$ for any $N$. Similar situations happen for $N_{P}[\rho_{S}^{\textrm{fact}}(0)]/(gt)^{2}$ when $N>1$ as seen in Eq. (54) and Fig. 6 (right column). Besides, as $N$ increases, $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/(gt)^{2}$ grows more and more fast with $|\rho_{eg}|$, so does $N_{P}[\rho_{S}^{\textrm{fact}}(0)]/(gt)^{2}$. Thus we infer that for states with large $N$, the cavity photon number $N_{P}$ are dominated by $|\rho_{eg}|$, and the memory effects are dominated by $N_{P}$. In Fig. 6(c) and (e), it is found that the states around $\rho_{eg}=0$ and $\rho_{ee}=0.5$ tend to take lower values of memory effects compared with their neighboring states, especially when $N$ is large. This might be intuitively interpreted by the fact that $\rho_{eg}=0$ and $\rho_{ee}=0.5$ leads to a maximally mixed state of the $N$-atom system. Besides, Fig. 6 shows that the (change of) the cavity photon number is not the only factor to determine the value of the memory effects. Other factors might include the (change of) coherence of the cavity state between the basis $|n\rangle$.
Figure 7: The degree of superradiance $S$ (left column) and the degree of memory-effect-enhancement $N_{M}/N_{M}^{\textrm{ind}}$ (from independent radiation to collective radiation, right column) for $\rho_{S}^{\textrm{fact}}(0)$ as functions of $\rho_{ee}$ and $|\rho_{eg}|$ in the early-time regime. The values of $S$ and $N_{M}/N_{M}^{\textrm{ind}}$ for $\rho_{ee}=0$ are not defined and not shown. The atom numbers are $N=2$, $3$, and $12$ from top to bottom. It is observed that large values of $S$ and $N_{M}/N_{M}^{\textrm{ind}}$ happen for low-excitation states with high coherence.
According to Eqs. (35) and (54), the degree of superradiance for $\rho_{S}^{\textrm{fact}}(0)$ is given by
$S[\rho_{S}^{\textrm{fact}}(0)]$ $=$ $\frac{N_{P}[\rho_{S}^{\textrm{fact}}(0)]}{\sum_{n=1}^{N}N_{P}^{(n)}(\rho_{S_{n}})}$ (55)
@@
$=$ $1+(N-1)\frac{|\rho_{eg}|^{2}}{\rho_{ee}}$
for state with $\rho_{ee}>0$. The state $\rho_{S}^{\textrm{fact}}(0)$ is superradiant when $|\rho_{eg}|$ is nonzero and $N>1$. Besides, the degree of superradiance grow quadratically with the single-atom coherence $|\rho_{eg}|$ for $N>1$. The maximum of $S[\rho_{S}^{\textrm{fact}}(0)]$ is $N$ when $\rho_{ee}\rightarrow 0$ and $|\rho_{eg}|=\sqrt{(1-\rho_{ee})\rho_{ee}}$ (its maximum under the semi-positivity constraint). We illustrate the degrees of superradiance for different atom numbers in Fig. 7 (left column). As done in last section, we also calculate the degree of memory-effect-enhancement $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/N_{M}^{\textrm{ind}}[\rho_{S}^{\textrm{fact}}(0)]$ in the early-time regime with Eqs. (29) and (30) for different atom numbers. These results are shown in Fig. 7 (right column). Unlike the case for the Dicke states, the dependence of $N_{M}/N_{M}^{\textrm{ind}}$ on $\rho_{ee}$ and $\rho_{eg}$ is not the same as that of $S$. But there are many similarities between them. When there are only one atom ($N=1$), $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/N_{M}^{\textrm{ind}}[\rho_{S}^{\textrm{fact}}(0)]=S[\rho_{S}^{\textrm{fact}}(0)]=1$ according to our definitions. For $N=2$, $S$ and $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/N_{M}^{\textrm{ind}}[\rho_{S}^{\textrm{fact}}(0)]$ are slightly different, implying that the memory effects are not merely determined by the cavity photon numbers for the factorized identical states. As the atom number increases (for $N\geqslant 3$), the surface shapes in the right-side figures become smoother and more similar to those in the left column, especially when $S$ is large. As shown in Fig. 7, strong memory-effect-enhancement happens for low-excitation states with high coherence (as $S$ does). Similar to Eq. (55), numerical simulations demonstrate that the maximum of $N_{M}[\rho_{S}^{\textrm{fact}}(0)]/N_{M}^{\textrm{ind}}[\rho_{S}^{\textrm{fact}}(0)]$ is also $N$ when $\rho_{ee}\rightarrow 0$ and $|\rho_{eg}|=\sqrt{(1-\rho_{ee})\rho_{ee}}$. These results imply that for states with high degree of superradiance, the (enhancement of) memory effects are dominated by the cavity photon number. Besides, numerical simulations show that for initial states whose spontaneous radiation can be enhanced by a common environment ($S>1$), their memory effects can always be enhanced simultaneously ($N_{M}[\rho_{S}^{\textrm{fact}}(0)]/N_{M}^{\textrm{ind}}[\rho_{S}^{\textrm{fact}}(0)]>1$). Therefore, we infer that the memory effects also contribute to the establishment of the superradiance for the factorized identical initial states.
@@ IV.3 Other initial states
Entanglement and coherence of the initial states may play important roles in superradiance [44, 24, 25, 26, 27], especially for the Dicke states. Next, we examine the influence of the coherence (or the entanglement among the atoms) in the dephased Dicke states $\rho^{\textrm{deph}}_{S}(0)$ on the memory effects as well as the superradiance. The initial state we consider is given by
$\rho^{\textrm{deph}}_{S}(0)=\lambda|JM\rangle\langle JM|+(1-\lambda)\mathcal{D}(|JM\rangle\langle JM|)$ (56)
where $0\leqslant\lambda\leqslant 1$. Here $\mathcal{D}$ turns $|JM\rangle\langle JM|$ into a fully dephased state with vanishing off-diagonal elements in the basis $|e(g),e(g),\cdots,e(g)\rangle$. Meanwhile, $\mathcal{D}(|JM\rangle\langle JM|)$ is a mixed state of separable states whose entanglement is zero. Thus $\lambda$ reflects the strength of entanglement or the value of coherence in $\rho^{\textrm{deph}}_{S}(0)$. With the help of numerical solutions, we find by mathematical induction that $\frac{N_{M}[\mathcal{D}(|JM\rangle\langle JM|)]}{(gt)^{2}}=\frac{1}{2}(J+M)$ and
$\frac{N_{M}[\rho^{\textrm{deph}}_{S}(0)]}{(gt)^{2}}$ $=$ $\lambda\frac{N_{M}(|JM\rangle\langle JM|)}{(gt)^{2}}$ (57)
@@
$\frac{N_{P}[\rho^{\textrm{deph}}_{S}(0)]}{(gt)^{2}}=$ $=$ $\lambda\frac{N_{P}(|JM\rangle)}{(gt)^{2}}$ (58)
@@
Similarly, $N_{M}[\rho^{\textrm{deph}}_{S}(0)]=\frac{1}{2}N_{P}[\rho^{\textrm{deph}}_{S}(0)]=\max_{\tau_{10},\tau_{21}}\Delta N_{ex}^{\textrm{atom}}[\rho^{\textrm{deph}}_{S}(0)]$ is satisfied for the dephased Dicke states in the early-time regime.
The degree of superradiance for the dephased Dicke states can be calculated from Eq. (58) by
$S[\rho^{\textrm{deph}}_{S}(0)]$ $=$ $\frac{N_{P}[\rho^{\textrm{deph}}_{S}(0)]}{\sum_{n=1}^{J+M}N_{P}(|e\rangle)}$ (59)
@@
for $J+M>0$. It is observed that the entanglement or coherence (represented by $\lambda$) in the depahsed Dicke state is necessary for superradiance. The degree of superradiance increases linearly with $\lambda$ for superradiant Dicke states. Next we discuss the degree of memory-effect-enhancement caused by a common environment. When the initial state is $\mathcal{D}(|JM\rangle\langle JM|)$ (fully dephased), we find that $N_{M}^{\textrm{ind}}[\mathcal{D}(|JM\rangle\langle JM|]/(gt)^{2}=N_{M}^{\textrm{ind}}[|JM\rangle\langle JM|]/(gt)^{2}=\frac{1}{2}(J+M)$. Moreover, for a partially dephased Dicke state described by Eq. (56), there is $N_{M}^{\textrm{ind}}[\rho^{\textrm{deph}}_{S}(0)]/(gt)^{2}=\frac{1}{2}(J+M)$ regardless of $\lambda$. Therefore, the degree of memory-effect-enhancement for $\rho^{\textrm{deph}}_{S}(0)$ is given by
$\frac{N_{M}[\rho^{\textrm{deph}}_{S}(0)]}{N_{M}^{\textrm{ind}}[\rho^{\textrm{deph}}_{S}(0)]}=(J-M)\lambda+1=S[\rho^{\textrm{deph}}_{S}(0)]$ (60)
as the case in Sec.IV A.
At the end of this section, we calculate the value of the normalized value of memory effects of a probabilistic mixture of Dicke states $\rho_{S}^{\textrm{mix}}(0)=\sum_{M}p_{M}|JM\rangle\langle JM|$ where $J=N/2$, $M=-J,-J+1,\cdots,J$, and $\sum_{M}p_{M}=1$. Using similar treatments as done in Eq. (45), we find that
$\frac{N_{M}[\rho_{S}^{\textrm{mix}}(0)]}{(gt)^{2}}\!$ $=$ $\!\frac{1}{4}[\!\sum_{M=-J}^{J-1}\!|p_{M+1}f(M+1)-p_{M}f(M)|$ (61)